OpenStudy (anonymous):

Please help me? I am so lost! The circle given by x^2+y^2-2y-11=0 can be written in standard form like this: x^2 + (y-k)^2=12. What is the value of k in this equation?

OpenStudy (anonymous):

expand out x^2 + y^2 -2yk +k^2 - 12 = 0 compare with the other one 2yk = 2y -11 = k^2 -12 work out k pssstt (try k = 1)

OpenStudy (schrodingers_cat):

Another way is just to complete the square for y x^2 + (y^2 -2y +1) -1 -11 =0

OpenStudy (anonymous):

where did the -2yk +k^2 - 12 come from?

OpenStudy (schrodingers_cat):

Then it simplifies down to x^2 + (y-1)^2 = 12

OpenStudy (anonymous):

I don't even know how you got the original one

OpenStudy (schrodingers_cat):

All you are doing is completing the square for y so it simplifies.

OpenStudy (anonymous):

i dont know how to complete a square...

OpenStudy (schrodingers_cat):

Completing the square is a very useful tool the comes handy in many situations and is most definitely worth learning!!!!!!!!!!!! Here is a link to a website that goes over the basics instead of me having to type it out http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (schrodingers_cat):

I hope this helps :)

OpenStudy (anonymous):

thanks